The barometric pressure at sea-level is 760mm of mercury while that on a mountain top is 735mm.If the density of air is assumed constant as 1.2kg/m3, what is the elevation of the mountain top.

Grade:Upto college level

1 Answers

askIITianexpert IITDelhi
14 years ago

Applying Bernoullie's theorem, 

               ΔP=ρairgh

      or, ΔP=(760-735)mm of mercury=25mm of mercury=(25mm)*(density of mercury)*g=(25*10-3)*(13.6*103)g

                                                                                                                                                            where,  ρmercury=13.6*103kg/m3

                                                                                                                                25mm=25*10-3m

       Hence height of mountain

                                         h=25*13.6/1.2=283.34m